OpenStudy (mrjackdzi):

is the derivative of e^x^2 = x^2 or 2x

OpenStudy (sweetburger):

chain rule is pertinent here

OpenStudy (coconutjj):

dy/dx = dy/du * du/dx

OpenStudy (coconutjj):

u = x^2 y = e^u dy/du = ? dy/dx = ?

OpenStudy (mrjackdzi):

yes i know that so it should be 2x * e^(x²) ?

OpenStudy (sweetburger):

yes

OpenStudy (sweetburger):

you are correct

OpenStudy (mrjackdzi):

Yes my home teaching is correct!

OpenStudy (mrjackdzi):

Thanks

OpenStudy (sweetburger):

np